# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v6

@@ 10.2 Theorem Statement
We have set up the containers by a GLOVER data structure. There are no back arcs, else we have a degree-doubling node. Next is lemma 7.2 (Interior Degree Doubling), which states that in every container we will have a node whose interior degree doubles.
@@
11:         Since there is no back arc, the container $C_{i+1}$ will be bound in size to be 1 less than the size of container $C_{i}$.
@@
Start with graph size $n$Select minimum degree node $v_{0}$ivide into containers $C_{1},C_{2},\ldots$Identify node $x_{i}$ in $C_{i}$ whose interior degree doublesMove to neighboring container $C_{i+1}$f back arc to prior containerLocate child of $x_{i}$ doubling $x_{i}$’s degreeHaltElse bound $|C_{i+1}|<|C_{i}|$Repeat from step 4
@@
Degree Doubling by Container Density
In an oriented graph $G$ with minimum degree node $v_{0}$ and containers $C_{1},\allowbreak C_{2},\allowbreak\ldots\allowbreak C_{n}$, as the containers shrink in size as their distance from the minimum degree node increases, and as the required interior degrees rise, a collision is inevitable.
@@
  • Case 1: The containers continue to shrink in size. Simultaneously, the interior degrees continue to grow in size. We reach a point where no feasible oriented graph is possible to accommodate the required set of nodes with interior degrees.
@@
Corollary 10.1 (Degree Doubling by Container Density) captures the essence of shrinking container sizes and increased number of nodes. This was represented by the yin and yang effects early on. We see it come full circle here. Ultimately, the interior neighbors represented the number of interior nodes, while the exterior neighbors represented the size of the containers.
In small graphs like 4 (Minimum Out-Degree 3 with Neighbors 1), we saw that this collision did not occur. To understand the reasoning behind this, understand that the container sizes are decreasing at the rate of 1, and the interior degrees are increasing at a rate of 1. The problem is that the number of arcs inside an oriented graph is growing at a rate of $O(n^{2})$. So, for small graphs, we are just not going out far enough to catch up. For larger graphs, though, we do start to see this.
@@
Occurrence of Last Dense Container
Let $G$ be an oriented graph represented by a GLOVER with minimum degree $k$. Then there will be a container that reaches maximal density. Further, if that container is $C_{i}$, it will be $i$-out-regular.
@@
We begin by stating lemma 6.1 (Regular Interior Degrees). While these are called interior degrees, these are the out-degrees of the nodes in a container. This lemma says that every container must have regular out-degrees.
@@
Thus, we have a series of regular containers that are getting smaller as we get further from the minimum degree none. At the same time, these containers are expected to hold fewer nodes of higher degree, i.e., a graph of more arcs. This is progressively getting to a denser container. As Corollary 10.1 (Degree Doubling by Container Density), this cannot continue forever. There must be a last, most dense container.
The Seymour Second Neighborhood Conjecture is made more beautiful by the appearance of these orientations of out-regular graphs as containers get smaller. Not only has the conjecture been proven true, but with an inherent symmetry and order. There is a visual pattern emerging inside these containers that is not just beautiful to graph theorists, but there is a structure that is easily appealing to non-mathematicians.
@@
Multiple Degree Doubling Nodes Let $G$ be an oriented graph represented by the GLOVER structure, with minimum degree $k$, and let the most dense graph in container $C_{i}$. Then every node in container $C_{i}$ has its degree doubled in the representation.
The discovery of multiple degree-doubling nodes lifts the Seymour Second Neighborhood Conjecture from a mere question of existence to one in abundance. There were many who doubted that one even existed. Now there are many.
Second is the way we find it. It’s the presence of an oriented regular graph as these containers decrease in size. There’s an element of mathematical beauty there.
Then third we can look at the applications. We may be looking for the person who has the most influence, but this is saying that there may not be just one person; there may be a whole set of people, a tightly connected group.
@@
Once we have this value, we will place every node and every edge of $G$ into a preprocessing array ready to be partitioned. Each edge is examined, contributing $O(|E|)$ complexity, while nodes contribute $O(|V|)$. Each node will individually be processed based on its distance from the minimum degree node, along with the connected outgoing arcs. This will require $O(|V|)$ comparisons and $O(|V|+|E|)$ stores into containers. This means that the partitioning has a complexity of $O(|V|+|E|)$.
@@
First, remember that we proved in lemma 7.2 (Interior Degree Doubles) that every container has a node whose interior degree doubles. Secondly, we also note that the lemma 6.1 (Regular Interior Degrees) implies that all the nodes are regular inside a container. Third, lemma 7.1 (Indistinguishable Regular Interior Nodes) lets us know that nodes within a container are indistinguishable from one another. The combination of these three lemmas let us know that every node inside a container has their interior degree double. This gives the complexity of finding a node whose interior degree has doubled $O(1)$.
What remains is to consider the complexity necessary to combine these interior degree nodes to solve the overall problem. Solving the full problem requires determining the number of containers necessary for this partition, which is bounded above by $O(|V|)$. The total number of containers cannot exceed the number of nodes, $|V|$, since containers are based on partitions. Processing each container’s exterior arcs and remaining possible nodes is bounded by $O(|V|+|E|)$. This gives the run time a complexity of $O(|V|+|E|)$.
@@ 11 Conclusion and Future Work
In this work, we explored the Seymour Second Neighborhood Conjecture in oriented graphs. We did so through the lenses of graph theory, combinatorics, algorithms, and data structures. This research helped us to use many mathematical tools like both total orders, strong and weak induction, recurrence relations, and the divide-and-conquer paradigm.
One of the key insights of this paper was the development of an algorithm that identifies nodes whose degree must double in any oriented graph. This effectively solves the Seymour Second Neighborhood Conjecture with an answer in the affirmative. The interesting thing about this algorithm was that it did not require any assumptions about the oriented graph or where the degree-doubling node was in the graph. Instead, the algorithm selects a minimum degree node and is able to partition the graph into containers and form a decreasing sequence of exteriors that is eventually bound to lead us to the solution.
This approach introduced a new perspective on the Seymour Second Neighborhood Conjecture. It was no longer framed as merely a graph theory problem. By utilizing the divide-and-conquer paradigm and solving different parts independently, this algorithm shows that the Seymour Second Neighborhood Conjecture could effectively be treated with computer science techniques. This suggests that there may be other open problems in mathematics that are not being given enough attention by computer scientists. Alternatively, it may be that there is not enough interdisciplinary communication about these problems and the potential techniques to solve them.
The Seymour Second Neighborhood Conjecture started as a question in 1990. Over the next 34 years it has inspired hundreds of mathematicians into graph theory, research, and some into computers. It has been talked about in books. on message boards, in classrooms, and at conferences. This problem has a long history.
With a solution now in hand, it’s time to begin looking forward. The Seymour Second Neighborhood Conjecture can be represented as a social media problem instead of an oriented graph. Here, the nodes are people, where no one follows back. Now instead of wondering if there is a person whose sphere of influence doubles, we would simply be searching for that person. We just need to implement the algorithm to find the person. We could have the graph represent epidemiology. Here the nodes will be infected people, and the arcs would be one-way infections. The question of interest would be, can we implement the algorithm to find the person who is a super-spreader and do so quickly? There are many more real-life examples where we can apply this problem and this algorithm. There are fields like security, telecommunications, urban planning, and neural networks. There are too many to name in this short paper.
@@ A Graph Level Order (GLOVER) on set $S$ can be defined as follows:
Let $v_{0}$ be a minimum degree node, $x$ be a node in $G$, and $y\in N^{+}(x)$ be in the component $C_{i}$. A back arc is defined as an arc $(y,z)$ such that $z\in N^{+}(y)$ and $x\in C_{j}$, where $j<i$.
Definition: Let $x\in G$ and $y,u\in N^{+}(x)$ and $(y,u)\in G$. Then $x$, $y$, $u$ form a transitive triangle.
@@ 14 Appendix 2: Applicability of the Graph Level Order
The deciding factor that held the algorithm together and helped solve the Seymour Second Neighborhood Conjecture was the Graph Level Order (GLOVER) data structure. This research did not begin as an investigation into data structures. Instead, a series of operations were conducted on the oriented graphs - partitioning, ordering those partitions, adding interior and exterior arcs, until what was left was an ordering of both the partition and nodes within the partition. Further, these interior and exterior arcs allow us to define relationships based on two different metrics, an outer metric and an inner metric. Unlike the Seymour Second Neighborhood Conjecture, exterior arcs do not always have to only be defined to the next container. That was a problem specific definition.
Moving forward, the GLOVER data structure should be well suited to tackle problems in graph theory and beyond. By introducing containers, another way of doing induction on graphs has been introduced. The concept of divide-and-conquer, which has dominated much of computer science, shows up in this paper as well. We are able to treat these containers independent of one another and solve their problems locally before bringing them back to the global problem. This shows how versatile the GLOVER data structure is moving forward.
@@
Next these nodes are not just treated like clusters inside these containers, but like nodes in graph theory. The interior and exterior arcs allow for common graph theory algorithms and techniques like Prim’s algorithm, breadth first search, or centrality to be called. What is even more interesting is that because we have identified the interior and exterior arcs already, we can choose to run these algorithms only on interior arcs or only on exterior arcs.
@@ 15 Appendix 3: Visualization and Accessibility of Oriented Graphs
When we take a course on graph theory, oriented graphs are probably not on the course load. Second neighbors or squares of graphs are probably not, either. To fully grasp this problem, though, we need to be able to delve into these definitions through visualizations.
For example, research on the Seymour Second Neighborhood Conjecture led this paper into the terms decreasing sequence property, induced subgraphs, containers, interior neighbors, exterior neighbors, and back arcs. The paper addressed how these terms were intertwined with each other. An example is where the paper says that there is a monotonically decreasing sequence.
James Robert Brown’s book Philosophy of Mathematics speaks of the importance of visualizations in mathematics. Proofs are still the foundation, but sometimes visualizations can serve as inspiration or a spark for those proofs.
@@
Readers are encouraged to explore the website for additional information and examples at: https://glovermethod.com This is an evolving project. Additional examples and applications will continue to be added. Readers are invited to explore the website. Please provide feedback on any aspects of the site or content for improvement.
Our goal here was to help users connect with the Seymour Second Neighborhood Conjecture. Hopefully, this tool not only serves as support for this paper, but as an educational resource in the fields of graph theory and computer science. Thank you for your engagement

### v7

@@ 10.2 Theorem Statement
We have set up the containers using a GLOVER data structure. There are no back arcs; else we will have a degree-doubling node. Next is lemma 7.2 (Interior Degree Doubling), which states that in every container we will have a node whose interior degree doubles.
@@
11:         Since there is no back arc, the container $C_{i+1}$ will be limited in size to be 1 less than the size of container $C_{i}$.
@@
Start with graph size $n$Select a minimum degree node $v_{0}$ivide into containers $C_{1},C_{2},\ldots C_{k}$dentify node $x_{i}$ in $C_{i}$ whose interior degree doublesMove to neighboring container $C_{i+1}$f back arc to a prior containerLocate the child of $x_{i}$ doubling $x_{i}$’s degreeHaltElse bound $|C_{i+1}|<|C_{i}|$Repeat from step 4
@@
Degree Doubling by Container Density In an oriented graph $G$ with a minimum degree node $v_{0}$ and containers $C_{1},\allowbreak C_{2},\allowbreak\ldots\allowbreak C_{n}$, as the containers shrink in size as their distance from the minimum degree node increases, and as the required interior degrees rise, a collision is inevitable.
@@
  • Case 1: The containers continue to shrink in size. Simultaneously, the interior degrees continue to grow. We reach a point where no feasible oriented graph is possible to accommodate the required set of nodes with interior degrees.
@@
Corollary 10.1 (Degree Doubling by Container Density) captures the essence of shrinking container sizes and an increased number of nodes. This was represented by the yin and yang effects early on. We see it come full circle here. Ultimately, the interior neighbors represented the number of interior nodes, while the exterior neighbors represented the size of the containers.
In small graphs like 4.3 (Minimum Out-Degree 3 with Neighbors 1), we saw that this collision did not occur. To understand the reasoning behind this, one must understand that the container sizes are decreasing at the rate of 1, and the interior degrees are increasing at a rate of 1. The problem is that the number of arcs inside an oriented graph is growing at a rate of $O(n^{2})$. So, for small graphs, we are just not going out far enough to catch up. For larger graphs, though, we do start to see this.
@@
Occurrence of Last Dense Container Let $G$ be an oriented graph represented by a GLOVER with minimum degree $k$. Then there will be a container that reaches maximal density. Further, if that container is $C_{i}$, it will be $i$-out-regular.
@@
We begin by stating lemma 6.1 (Regular Interior Degrees). While these are called interior degrees, these are the out-degrees of the nodes in a container. This lemma indicates that every container must have regular out-degrees.
@@
Thus, we have a series of regular containers that are getting smaller as we get further from the minimum degree none. At the same time, these containers are expected to hold fewer nodes of higher degree, i.e., a graph of more arcs. This is progressively getting to a denser container. As described in corollary 10.1 (Degree Doubling by Container Density), this cannot continue forever. There must be a last, most dense container. ∎
The SSNC is made more aesthetic by the appearance of these orientations of out-regular graphs as containers get smaller. Not only has the conjecture been proven true, but with an inherent symmetry and order. There is a visual pattern emerging inside these containers that is not just important to graph theorists but also possesses a structure that is easily appealing to non-mathematicians.
@@
Multiple Degree Doubling Nodes Let $G$ be an oriented graph represented by the GLOVER structure, with minimum degree $\delta$, and let the most dense graph be in container $C_{i}$. Then every node in container $C_{i}$ has its degree doubled in the representation.
The discovery of multiple degree-doubling nodes elevates the SSNC from a mere question of existence to one in abundance. There were many who doubted that one even existed. Now there are many.
Second is the way we find it. It’s the presence of a last dense container with an oriented regular graph as these containers decrease in size. We see this occur as the interior degrees are forced to increase as the distance from $v_{0}$ is increased. These interior degrees are set up to prevent a parent node’s degree from doubling. Ultimately though, they are unable to do that, as we see the collision of the container shrinkage and interior degree requirements.
Then third we can look at the applications. We may be looking for the person who has the most influence, but this is saying that there may not be just one person; there may be a whole set of people, a tightly connected group. Dense regular graphs have many applications in themselves, from areas like expander graphs to network topology design to coding theory.
@@
Once we have this value, we will place every node and every edge of $G$ into a pre-processing array ready to be partitioned. Each edge is examined, contributing $O(|E|)$ complexity, while nodes contribute $O(|V|)$. Each node will individually be processed based on its distance from the minimum degree node, along with the connected outgoing arcs. This will require $O(|V|)$ comparisons and $O(|V|+|E|)$ stores into containers. This means that the partitioning has a complexity of $O(|V|+|E|)$.
@@
First, remember that we proved in lemma 7.2 (Interior Degree Doubles) that every container has a node whose interior degree doubles. Secondly, we also note that the lemma 6.1 (Regular Interior Degrees) implies that all the nodes are regular inside a container. Third, lemma 7.1 (Indistinguishable Regular Interior Nodes) lets us know that nodes within a container are indistinguishable from one another. The combination of these three lemmas lets us know that every node inside a container has their interior degree doubled. This gives the complexity of finding a node whose interior degree has doubled $O(1)$.
What remains is to consider the complexity necessary to combine these interior degree nodes to solve the overall problem. Solving the full problem requires determining the number of containers necessary for this partition, which is bounded above by $O(|V|)$. The total number of containers cannot exceed the number of nodes, $|V|$, since containers are based on partitions. Processing each container’s exterior arcs and remaining possible nodes is bounded by $O(|V|+|E|)$. This gives the run time a complexity of $O(|V|+|E|)$. ∎
@@
We opened by showing several applications of the SSNC, including social media, social network modeling and network analysis. This highlights the necessity for an algorithm that can work efficiently in practice. This proof that the algorithm has a complexity of $O(|V|+|E|)$, where $|V|$ is the number of vertices and $|E|$ is the number of edges in the graph is important. This algorithm’s linear time complexity shows that even enormous graphs may be processed effectively. That makes it potentially applicable to real-world networks with millions or billions of nodes and edges. This suggests that the algorithm is not only theoretically sound but also practically feasible for implementation and use in real-world scenarios, such as analyzing social networks, identifying influential nodes in communication networks, or understanding the spread of information.
@@ 11 Conclusion and Future Work
In this work, we explored the Seymour Second Neighborhood Conjecture (SSNC) in oriented graphs. We did so through the lenses of graph theory, combinatorics, algorithms, and data structures. This research helped us to use many mathematical tools like both total orders, strong and weak induction, recurrence relations, and the divide-and-conquer paradigm.
One of the key insights of this paper was the development of an algorithm that identifies nodes whose degree must double in any oriented graph. This effectively solves the SSNC with an answer in the affirmative. The interesting thing about this algorithm was that it did not require any assumptions about the oriented graph or where the degree-doubling node was in the graph. Instead, the algorithm selects a minimum degree node and can partition the graph into containers and form a decreasing sequence of exteriors that is eventually bound to lead us to the solution.
This approach introduced a new perspective on the SSNC. It was no longer framed as merely a graph theory problem. By utilizing the divide-and-conquer paradigm and solving different parts independently, this algorithm shows that the SSNC could effectively be treated with computer science techniques. This suggests that there may be other open problems in mathematics that are not being given enough attention by computer scientists. Alternatively, it may be that there is not enough interdisciplinary communication about these problems and the potential techniques to solve them.
The SSNC started as a question in 1990. Over the next 34 years it has inspired hundreds of mathematicians into graph theory, research, and some into computers. It has been talked about in books. on message boards, in classrooms, and at conferences. This problem has a long history.
With a solution now in hand, it’s time to begin looking forward. The SSNC can be represented as a social media problem instead of an oriented graph. Here, the nodes are people, where no one follows back. Now instead of wondering if there is a person whose sphere of influence doubles, we would simply be searching for that person. We just need to implement the algorithm to find the person. We could have the graph represent epidemiology. Here the nodes will be infected people, and the arcs would be one-way infections. The question of interest would be, can we implement the algorithm to find the person who is a super-spreader and do so quickly? There are many more real-life examples where we can apply this problem and this algorithm. There are fields like security, telecommunications, urban planning, and neural networks. There are too many to name in this short paper.
We have presented a linear time algorithm that can work on large graphs in these problems. By constructing the algorithm for the SSNC and proving that the algorithm will find a degree-doubling node, we have proved that the algorithm is theoretically sound.
@@
Finally, this research offered an alternative method for the SSNC. We focused on identifying a node whose out-degree doubles in the graph’s square rather than just showing that it exists. To do this, we developed an effective algorithm and proved its complexity. Other problems in graph theory and related domains may benefit from this change in emphasis from existential proofs to algorithmic solutions. This may encourages researchers to explore constructive approaches with a focus on practical applications. Hopefully, this work will inspire further research on constructive approaches to other graph theory problems. Also, we encourage greater collaboration between mathematicians and computer scientists to explore the algorithmic and practical implications of theoretical results.
@@ A Graph Level Order (GLOVER) on set $S$ can be defined as follows:
Let $v_{0}$ be a minimum degree node, $x$ be a node in $G$, and $y\in N^{+}(x)$ be in the component $C_{i}$. A back arc is defined as an arc $(y,z)$ such that $z\in N^{+}(y)$ and $x\in C_{j}$, where $j<i$. Definition: Let $x\in G$ and $y,u\in N^{+}(x)$ and $(y,u)\in G$. Then $x$, $y$, $u$ form a transitive triangle.
@@ 14 Appendix 2: Applicability of the Graph Level Order
The deciding factor that held the algorithm together and helped solve the SSNC was the Graph Level Order (GLOVER) data structure. This research did not begin as an investigation into data structures. Instead, a series of operations were conducted on the oriented graphs—partitioning, ordering those partitions, and adding interior and exterior arcs—until what was left was an ordering of both the partition and nodes within the partition. Further, these interior and exterior arcs allow us to define relationships based on two different metrics, an outer metric and an inner metric. Unlike the SSNC, exterior arcs do not always have to only be defined to the next container. That was a problem-specific definition.
Moving forward, the GLOVER data structure should be well suited to tackle problems in graph theory and beyond. By introducing containers, another way of doing induction on graphs has been introduced. The concept of divide-and-conquer, which has dominated much of computer science, shows up in this paper as well. We can treat these containers independently of one another and solve their problems locally before bringing them back to the global problem. This shows how versatile the GLOVER data structure is moving forward.
@@
Next, these nodes are not just treated like clusters inside these containers but like nodes in graph theory. The interior and exterior arcs allow for common graph theory algorithms and techniques like Prim’s algorithm, breadth-first search, or centrality to be called. What is even more interesting is that because we have identified the interior and exterior arcs already, we can choose to run these algorithms only on interior arcs or only on exterior arcs.
@@ 15 Appendix 3: Visualization and Accessibility of Oriented Graphs
When we take a course on graph theory, oriented graphs are probably not on the course load. Second neighbors or squares of graphs are probably not either. To fully grasp this problem, though, we need to be able to delve into these definitions through visualizations.
For example, research on the SSNC led this paper into the terms decreasing sequence property, induced subgraphs, containers, interior neighbors, exterior neighbors, and back arcs. The paper addressed how these terms were intertwined with each other. An example is where the paper says that there is a monotonically decreasing sequence.
James Robert Brown’s book ”Philosophy of Mathematics” speaks of the importance of visualizations in mathematics. Proofs are still the foundation, but sometimes visualizations can serve as inspiration or a spark for those proofs.
@@
Readers are encouraged to explore the website for additional information and examples at https://glovermethod.com. This is an evolving project. Additional examples and applications will continue to be added. Readers are invited to explore the website. Please provide feedback on any aspects of the site or content for improvement. Our goal here was to help users connect with the SSNC. Hopefully, this tool not only serves as support for this paper but also as an educational resource in the fields of graph theory and computer science. Thank you for your engagement.